OpenStudy (anonymous):

given iken and ne=12 and ik =2x-4 what is the value of x

OpenStudy (anonymous):

\(ne=12\) and \(ik=2x-4\) so \(ilne =(2x-4)\times 12=24x-48\)

OpenStudy (anonymous):

so i guess you need to solve \(24x-48=12\)